OpenStudy (twizttiez):

Choose the correct answer. Allan and Dave bowl together and their combined total score for one game was 375 points. Allan’s score was 60 less than twice Dave’s. What were their scores? Which is a system of equations to model the problem if x represents Dave’s score and y represents Allan’s score? A. x + y = 60 y = 2x – 375 B. x + y = 375 y = 2x – 60 C. x + y = 375 y = 2x + 60 D. x – y = 375 y = 2x – 60

OpenStudy (bohotness):

you called me?

OpenStudy (twizttiez):

Yeah earlier i need help

OpenStudy (bohotness):

okaya

OpenStudy (bohotness):

what do you think the answer is?

OpenStudy (twizttiez):

I think this is B

OpenStudy (twizttiez):

x + y = 375 y = 2x - 60 x + 2x - 60 = 375 3x - 60 = 375 3x = 375 + 60 3x = 435 x = 435/3 x = 145 x + y = 375 145 + y = 375 y = 375 - 145 y = 230 Dave (x) scored 145 Allen (y) scored 230

OpenStudy (bohotness):

okay Since their added scores are 375, and one is x and the other is y, then you know x + y = 375. You can eliminate the 4th choice.

OpenStudy (bohotness):

We need to write this as an equation: "Allan’s score was 60 less than twice Dave’s. "

OpenStudy (twizttiez):

Ok

OpenStudy (bohotness):

Now we need to write "Allan’s score was 60 less than twice Dave’s. " as an equation. We are told Allan's score is y and Dave's score is x. Let's do that below the statement. "Allan’s score was 60 less than twice Dave’s. " y's score was 60 less than twice x's y = 2x - 60

OpenStudy (twizttiez):

Ok

OpenStudy (twizttiez):

So i would do x + y = 375 y = 2x - 60 x + 2x - 60 = 375 3x - 60 = 375 3x = 375 + 60 3x = 435 x = 435/3 x = 145 x + y = 375 145 + y = 375 y = 375 - 145 y = 230 Dave (x) scored 145 Allen (y) scored 230

OpenStudy (bohotness):

corrrcet

OpenStudy (twizttiez):

Yay

OpenStudy (bohotness):

thank you for the medal

OpenStudy (twizttiez):

No prob have a nice evening

OpenStudy (bohotness):

you too